OpenStudy (anonymous):

Look at the pattern below. 52, 104, 208, 416, ... Which number sentence can be used to determine n, the 8th number in the pattern?

OpenStudy (amistre64):

divide it all by 52, what do we get?

OpenStudy (anonymous):

ummm? a n = 26 ÷ 8 b n = 26 × 82 c n = 26 × 28 d n = (26 × 8) - 4 i got B as my answer

OpenStudy (amistre64):

52(1,2,4,8,...) a_n+1 = 52 * 2an an = 52 * 2^n is what it looks like to me

OpenStudy (amistre64):

if n starts at 0 :) so n=7 is the 8th term